the length is 14 cm, width is 8 cm. what is the expression for the perimeter of a rectangle?
Found 2 solutions by solver91311, drj:
Answer by solver91311(24713)



The perimeter, , of a rectangle is given by:



where is the length and is the width.

John


Answer by drj(1380)

The length is 14 cm, width is 8 cm. what is the expression for the perimeter of a rectangle?

Step 1. Perimeter P means that we add up all four sides of the rectangle.

Step 2. Let W be the width and L be the length.

Step 3. Then P=W+W+L+L=2W+2L=2(W+L)

Step 4. Given L=14 cm and W=8 cm, then P=2(14+8)=44.

Step 5. ANSWER: The perimeter is 44 cm where perimeter P is found as P=2(W+L).

I hope the above steps were helpful.

For FREE Step-By-Step videos in Introduction to Algebra, please visit http://www.FreedomUniversity.TV/courses/IntroAlgebra and for Trigonometry visit http://www.FreedomUniversity.TV/courses/Trigonometry.

Good luck in your studies!

Respectfully,
Dr J